A biologist is studying a new hybrid tomato. It is known that the seeds of this hybrid tomato have probability of 0.70 of germinating. The biologist plants six seeds. What is the average number of seeds we would expect to germinate?

6

0.7

4.2

3
Answer #1

Solution :

Given that,

p = 0.70

n = 6

Using binomial distribution,

Mean = = n * p = 6 * 0.70 = 4.2

correct option is = 4.2